In a lot of Tv manufactured,each tv checked and found Avg.defect per tv =4,w hat is the possibilities of maximum defect in any tv?


1
Expert's answer
2021-12-22T14:57:34-0500

using Poisson distribution, probability that there are no defects in tv:

"P(x=0)=e^{-\lambda}=e^{-4}=0.0183"

then probability that there are defects in tv:

"P(x\ge0)=1-0.0183=0.9817"

mean:

"\mu=\lambda=4"

standard deviation:

"\sigma=\sqrt{\lambda}=2"


for normal distribution:

for "P=0.0183" :

"z=\frac{x-\mu}{\sigma}=\frac{x-4}{2}=2.09"


possibilities of maximum defect in any tv:

"x=2\cdot2.09+4=8.18\approx8" defects